OpenStudy (anonymous):

f(x) = a^2 - 1 , g(x) = a/(a + 1) .. Can someone help me figure out the f+g , f-g , f*g and f/g

OpenStudy (anonymous):

\[f(x)=x ^{2}-1\] \[g(x)=\frac{ x }{ x+1 }\] To solive f+g, simply add both equations \[f+g=\left( x^2+1 \right)+\left( \frac{ x }{ x+1 } \right)\]

OpenStudy (anonymous):

I don't think there is any need to simplify, unless it states in the quesiton

OpenStudy (anonymous):

use this idea to solve the other new functions